# Controlling discrete semilinear wave equations toward flocks

### v1

@@ 4. Connection to optimal control and Hamilton–Jacobi equations
The ODE system (1.3) is clearly nonlocal in nature. It can be formulated in the optimal control framework as follows.
Write $X_{1}=(x_{l})_{l\in\overline{\Lambda}}$ and $X_{2}=(v_{l})_{l\in\overline{\Lambda}}$ as vectors in $\mathbb{R}^{D^{n}}$. Denote by $X=(X_{1},X_{2})\in\mathbb{R}^{2D^{n}}$. Then, (1.3) can be written as
@@
with given initial condition
@@
In this framework, $W$ is the control and the admissible set is
@@
Here, $|\cdot|_{1}$ is the usual $1$-norm in $\mathbb{R}^{D^{n}}$. Therefore, $U$ satisfies the following eikonal type equation in the viscosity sense
@@
See [13, Chapter 2]. We note that this is a very complicated PDE posed in high dimensions as the vector field $b$ contains much information about the dynamics. The value function $U(X)$ gives us the minimal time to reach a moving flock starting from $X\in\mathbb{R}^{2D^{n}}$, but does not give us directly an optimal strategy to reach a moving flock. By Theorem 1.1, we have that $U$ is finite. As the Hamiltonian $H$ here is noncoercive in $P$, the regularity of the viscosity solution $U$ is not yet known.

### v3

@@ 4. Connection to optimal control and Hamilton–Jacobi equations
The goal of the control strategy above is to drive the system to a flock in a finite amount of time. Although the strategy strives to do this efficiently, it is by no means optimal. A related problem is to find the minimal time it takes to start from a given state and reach a flock with a control that has magnitude less than $M$. There is a rich theory behind this kind of optimal control problem. Here, we connect this optimal problem to a Hamilton-Jacobi problem.
The first step is to consider an ODE that does not have an interaction term like the one in (1.3). Thus, we write $X_{1}=(x_{l})_{l\in\overline{\Lambda}}$ and $X_{2}=(v_{l})_{l\in\overline{\Lambda}}$ as vectors in $\mathbb{R}^{D^{n}}$. Denote by $X=(X_{1},X_{2})\in\mathbb{R}^{2D^{n}}$. Then, (1.3) can be written as
@@
with the given initial condition
@@
In this framework, $W$ is the control, and the admissible set is
@@
Here, $|\cdot|_{1}$ is the usual $1$-norm in $\mathbb{R}^{D^{n}}$. Therefore, $U$ satisfies the following eikonal-type equation in the viscosity sense
@@
See [32, Chapter 2]. We note that this is a very complicated PDE posed in high dimensions, as the vector field $b$ contains much information about the dynamics. The value function $U(X)$ gives us the minimal time to reach a moving flock starting from $X\in\mathbb{R}^{2D^{n}}$, but does not give us directly an optimal strategy to reach a moving flock. By Theorem 1.1, we have that $U$ is finite. As the Hamiltonian $H$ here is noncoercive in $P$, the regularity of the viscosity solution $U$ is not yet known.
@@
5. Conclusion
This work is the first in controlling Klein-Gordon chains and lattices. Due to the complexity of the nonlinear ODE system under consideration, we focus here on a periodic domain and on controls applied at every node.
Our analysis provides insight into the system’s dynamics. Building on this understanding, we plan to investigate the problem with sparse controls, where the evolution of neighboring nodes will be used to influence nodes for which $u_{l}\equiv 0$.
In future work, we also aim to study the problem without assuming periodicity. This leads to an infinite-dimensional ODE system, for which the Lyapunov functional developed in this paper is not well defined. Our intended strategy is to first steer the infinite-dimensional system toward a periodic regime and then apply the results obtained here. Thus, the present paper constitutes a crucial first step in this broader research program.